# Addressing the Current Challenges of Quantum Machine Learning through Multi-Chip Ensembles

### v1

@@ Appendix G Experimental Model Design
Figure 3: Processing High Dimensional Data: Single-chip vs Multi-chip Ensemble VQCs
@@
Table 1: Experimental Model Design: Proof-of-Concept
@@ G.6 QCNN Architecture for PhysioNet EEG
We implement a quantum convolutional neural network (QCNN) [4] and adapt it to our multi-chip setting and the high-dimensional PhysioNet EEG input ($62\times 52=3264$ features).
@@ Appendix H Experimental Results: FashionMNIST, CIFAR-10, PhysioNet EEG
Figure 4: Experimental Results on FashionMNIST. DimReduc 2-Chip and 4-Chip indicate multi-chip ensemble VQC models with classical dimension reduction. Ensemble 98-Chip denotes multi-chip ensemble VQC model without classical dimension reduction.
@@
Figure 5: Experimental Results on CIFAR-10. DimReduc 2-Chip, 3-Chip, 4-Chip, and 6-Chip indicate multi-chip ensemble VQC models with classical dimension reduction. Ensemble 256-Chip denotes multi-chip ensemble VQC model without classical dimension reduction.
Figure 6: Model Performance
Figure 7: Experimental Results on PhysioNet EEG dataset. Ensemble 272-Chip denotes multi-chip ensemble VQC model without classical dimension reduction.

### v2

@@ D.3 Summary
D.4 Empirical Verification of Barren Plateau Mitigation
To verify our theoretical results empirically, we measured both the entangling capability and gradient variance across different VQC architectures with a fixed total of 8 qubits. Table 1 demonstrates the inverse relationship between entanglement and gradient variance. As the number of chips increases from single-chip to 4-chip ensembles, the entangling capability decreases while gradient variance increases by nearly an order of magnitude. This confirms our theoretical prediction that partitioning circuits into multi-chip ensembles reduces global entanglement and consequently mitigates the barren plateau problem by increasing gradient variance.
Table 1: Level of Quantum Entanglement and Variance of Gradients in Single-Chip vs Multi-Chip Ensemble VQCs. Larger values of entangling capability indicate higher levels of quantum entanglement. Larger values of gradient variance indicate lower risk of barren plateaus.
VQC Model Entangling Capability Variance of Gradients
Single-Chip 3.06998 5.95013e-8
2-Chip Ensemble 2.86057 1.39196e-7
4-Chip Ensemble 2.82801 3.51195e-7
@@ Appendix G Experimental Model Design
Figure 3: Conceptual comparison of data processing pipelines for high-dimensional inputs. (a) Single-Chip VQC: High-dimensional data typically requires a classical encoder for dimension reduction before being processed by a single VQC, followed by a classical decoder. This can lead to information loss. (b) Multi-Chip Ensemble VQC (Reduced): Demonstrates a multi-chip ensemble approach where the data, after classical dimension reduction by an encoder, is partitioned and processed by multiple smaller VQCs, with outputs combined by a classical decoder. (c) Multi-Chip Ensemble VQC (Full-Dimensional): Our proposed approach, where high-dimensional input data is directly partitioned (e.g., feature-wise) across multiple independent VQCs without prior classical dimension reduction. Outputs are combined by a classical decoder. This architecture aims to leverage the full dimensionality of the data while distributing the quantum computational load.
@@
Table 2: Experimental Model Design: Standard Benchmark Evaluation
@@ G.6 QCNN Architecture for PhysioNet EEG
We implement a quantum convolutional neural network (QCNN) [5] and adapt it to our multi-chip setting and the high-dimensional PhysioNet EEG input ($64\times 51=3264$ features).
@@ Appendix H Experimental Results: FashionMNIST, CIFAR-10, PhysioNet EEG
Figure 4: Experimental Results on FashionMNIST. (a) Model Performance: Training and validation loss (MSE) over epochs. (b) Generalizability: Generalization error. (c) Noise Resilience: Quantum error over epochs, including comparison with ZNE. Model naming conventions (’Ensemble 98-Chip’, ’DimReduc 2/4-Chip’, ’Classical’, ’Single-Chip’, ’ZNE Single-Chip’) are consistent with Figure 2. The results corroborate trends observed on MNIST, showing superior performance and robustness of the multi-chip ensemble approaches.
@@
Figure 5: Experimental Results on CIFAR-10. (a) Model Performance: Training and validation loss (MSE) over epochs. (b) Generalizability: Generalization error. (c) Noise Resilience: Quantum error over epochs, including comparison with ZNE. The Ensemble 256-Chip model processes full-dimensional CIFAR-10 data. ’DimReduc X-Chip’ models use X chips with prior classical dimension reduction. The results further validate the benefits of the multi-chip ensemble framework on a more complex image dataset.
Figure 6: Experimental Results on PhysioNet EEG dataset using QCNN. The plot shows the AUROC for training and validation sets over epochs. Compared are a classical CNN baseline, a single-chip QCNN (with dimension reduction), and our Ensemble 272-Chip QCNN which processes the full 3264-dimensional spatio-temporal features without classical dimension reduction. The multi-chip ensemble demonstrates superior performance and generalization (smaller train-validation gap) on this real-world, high-dimensional biomedical dataset.
Appendix I Code Availability
All analyses results in this study can be obtained using the codes shown in the link below:
https://github.com/JHPark9090/MultiChip-QML